HEBREW LETTER

type:  print: type: print:
\aleph ℵ \beth ℶ
\daleth ℸ \gimel ℷ

All symbols but \aleph need the amssymb package.

GREEK CHARACTERS

type:  print: type:  print: type:  print:
\alpha  α \beta  β \gamma  γ
\digamma  𝟋 \delta  δ \epsilon  ϵ
\varepsilon ε \zeta  ζ  \eta  η
\theta  θ \vartheta  ϑ \iota  ι
\kappa  κ \varkappa κ \lambda  λ
\mu  μ \nu  ν \xi  ξ
\pi  π \varpi  ϖ \rho  ρ
\varrho  ϱ \sigma  σ \varsigma ς
\tau  τ  \upsilon  υ \phi  ϕ
\varphi  φ \chi  χ \psi  ψ
\omega  ω

\digmma and \varkappa require the amssymb package.

type:  print: type:  print:
\Gamma Γ \varGamma Γ
\Delta Δ \varDelta Δ
\Lambda Λ \varLambda Λ
\Theta  Θ \varTheta  Θ
\Xi  Ξ \varXi  Ξ
\Pi  Π \varPi  Π
\Sigma  Σ \varSigma  Σ
\Upsilon  Υ \varUpsilon  Υ
\Phi  Φ \varPhi  Φ
\Psi  Ψ \varPsi  Ψ
\Omega  Ω \varOmega  Ω

All symbols whose name begins with var need the amssymb package.

L^{A}TEX BINARY RELATIONS

type:  print: type:  print:
\in  ∈ \ni  ∋
\leq  ≤ \geq  ≥
\ll  ≪ \gg  ≫
\prec  ≺ \succeq  ⪰
\sim  ∼ \cong  ≅
\simeq  ≃ \approx  ≈
\equiv  ≡ \doteq  ≐
\subset  ⊂ \supset  ⊃
\subseteq  ⊆ \supseteq  ⊇
\sqsubseteq ⊑ \sqsupseteq ⊒
\smile  ⌣ \frown  ⌢
\perp  ⊥ \models  ⊧
\mid  | \parallel  ∥
\vdash  ⊢ \dashv  ⊣
\propto  ∝ \asymp  ≍
\bowtie  ▷◁
\sqsubset ⊏ \sqsupset ⊐
\Join  ⋈

The latter three symbols need the latexsym package.

AMS BINARY RELATIONS

type:  print: type:  print:
\leqslant  ⩽ \geqslant  ⩾
\eqslantless  ⪕ \eqslantgtr  ⪖
\lesssim  ≲ \gtrsim  ≳
\lessapprox  ⪅ \gtrapprox  ⪆
\approxeq  ≊
\lessdot  ⋖ \gtrdot  ⋗
\lll  ≪ \ggg  ≫
\lessgtr  ≶ \gtrless  ≷
\lesseqgtr  ⋚ \gtreqless  ⋛
\lesseqqgtr  ⪋ \gtreqqless  ⪌
\doteqdot  ≑ \eqcirc  ≖
\circeq  ≗ \fallingdotseq  ≒
\risingdotseq  ≓ \triangleq  ≜
\backsim  ∽ \thicksim  ∼
\backsimeq  ⋍ \thickapprox  ≈
\preccurlyeq  ≼ \succcurlyeq  ≽
\curlyeqprec  ⋞ \curlyeqsucc  ⋟
\precsim  ≾ \succsim  ≿
\precapprox  ≾ \succapprox  ≿
\subseteqq  ⫅ \supseteqq  ⫆
\Subset  ⋐ \Supset  ⋑
\vartriangleleft  ◁ \vartriangleright  ▷
\trianglelefteq  ⊴ \trianglerighteq  ⊵
\vdash  ⊢ \Vdash  ⊩
\Vvdash  ⊪
\smallsmile  ⌣ \smallfrown  ⌢
\shortmid  ∣  \shortparallel  ∥
\bumpeq  ≏ \Bumpeq  ≎
\between  ≬ \pitchfork  ⋔
\varpropto  ∝ \backepsilon  ϶
\blacktriangleleft ◀ \blacktriangleright ▶
\therefore  ∴ \because  ∵

All symbols require the amssymb package.

AMS NEGATED BINARY RELATIONS

type:  print: type:  print:
\ne  ≠ \notin  ∉
\nless  ≮ \ngtr  ≯
\nleq  ≰ \ngeq  ≱
\nleqslant  ⩽̸ \ngeqslant  ⩾̸
\nleqq  ≦̸ \ngeqq  ≧̸
\lneq  ⪇ \gneq  ⪈
\ngeqq  ≧̸ \gneqq  ≩
\lvertneqq  ≨ \gvertneqq  ≩
\lnsim  ⋦ \gnsim  ⋧
\lnapprox  ⪉ \gnapprox  ⪊
\nprec  ⊀ \nsucc  ⊁
\npreceq  ⊀ \nsucceq  ⊁
\precneqq  ⪵ \succneqq  ⪶
\precnsim  ⋨ \succnsim  ⋩
\precnapprox  ⪹ \succnapprox  ⪺
\nsim  ≁ \ncong  ≇
\nshortmid  ∤  \nshortparallel  ∦
\nmid  ∤  \nparallel  ∦
\nvdash  ⊬ \nvDash  ⊭
\ntriangleleft  ⋪ \ntriangleright  ⋫
\ntrianglelefteq ⋬ \ntrianglerighteq ⋭
\nsubseteq  ⊈ \nsupseteq  ⊉
\nsubseteqq  ⫅̸ \nsupseteqq  ⫆̸
\subsetneq  ⊊ \supsetneq  ⊋
\varsubsetneq  ⊊ \varsupsetneq  ⊋
\subsetneqq  ⫋ \supsetneqq  ⫌
\varsubsetneqq ⫋ \varsupsetneqq  ⫌

All symbols but \ne require the amssymb package.

BINARY OPERATIONS

type:  print: type:  print:
\pm  ± \mp  ∓
\times  × \cdot  ·
\circ  ∘ \bigcirc  ◯
\div  ÷ \diamond  ⋄
\ast  ∗ \star  ⋆
\cap  ∩ \cup  ∪
\sqcap  ⊓ \sqcup  ⊔
\wedge  ∧ \vee  ∨
\triangleleft  ◁ \triangleright  ▷
\bigtriangleup △ \bigtriangledown ▽
\oplus  ⊕ \ominus  ⊖
\otimes  ⊗ \oslash  ⊘
\odot  ⊙ \bullet  •
\dagger  † \ddagger  ‡
\setminus  \ \uplus  ⊎
\wr  ≀  \amalg  ⨿
\lhd  ⊲ \rhd  ⊳
\unlhd  × \unrhd  ⊵
type:  print: type:  print:
\dotplus  ∔ \centerdot  ▪
\ltimes  ⋉ \rtimes  ⋊
\leftthreetimes ⋋ \rightthreetimes  ⋌
\circleddash ⊝ \smallsetminus ∖
\barwedge  ⊼ \doublebarwedge ⩞
\curlywedge  ⋏ \curlyvee  ⋎
\veebar  ⊻ \intercal  ⊺
\Cap  ⋒ \Cup  ⋓
\circledast  ⊛ \circledcirc  ⊚
\boxminus  ⊟ \boxtimes  ⊠
\boxdot  ⊡ \boxplus  ⊞
\divideontimes ⋇
\And  &

This table is divided into four parts. The first part contains the binary operations in L^{A}TEX. The second part requires the latexsym package. The third part contains the AMS additions; they require the amssymb package.

The symbol \And requires the amsmath package.

Arrows

type:  print: type:  print:
\leftarrow  ← \rightarrow or \to  →
\longleftarrow  ⟵ \longrightarrow  ⟶
\Leftarrow  ⇐ \Rightarrow  ⇒
\longleftarrow  ⟵ \longrightarrow  ⟶
\leftrightarrow  ↔ \longleftrightarrow ⟷
\Leftrightarrow  ⇔ \Longleftrightarrow ⟺
\uparrow  ↑ \downarrow  ↓
\Uparrow  ⇑ \Downarrow  ⇓
\updownarrow  ↕ \Updownarrow  ⇕
\nearrow  ↗ \searrow  ↘
\swarrow  ↙ \nwarrow  ↖
\mapsto  ↦ \longmapsto  ⟼
\hookleftarrow  ↩ \hookrightarrow  ↪
\leftharpoonup  ↼ \rightharpoonup  ⇀
\leftharpoondown ↽ \rightharpoondown ⇁
\rightleftharpoons ⇌
\leadsto  ⇝
type:  print: type:  print:
\leftleftarrows  ⇇ \rightrightarrows  ⇉
\leftrightarrows  ⇆ \rightleftarrows  ⇄
\Lleftarrow  ⇚ \Rrightarrow  ⇛
\twoheadleftarrow  ↞ \stwoheadrightarrow ↠
\leftarrowtail  ↢ \rightarrowtail  ↣
\looparrowleft  ↫ \looparrowright  ↬
\upuparrows  ⇈ \downdownarrows  ⇊
\upharpoonleft  ↿ \upharpoonright  ↾
\leftrightsquigarrow ↭ \rightsquigarrow  ⇝
\multimap  ⊸
\nleftarrow  ↚ \nrightarrow  ↛
\nLeftarrow  ⇍ \nRightarrow  ⇏
\nleftrightarrow  ↮ \nLeftrightarrow  ⇎

This table is divided into three parts. The top part contains the symbols provided by L^{A}TEX; the last command, \leadsto, requires the latexsym

package. The middle part table contains the AMS arrows; they all require the amssymb package. The bottom table lists the negated arrow symbols; they

also require amssymb.

Miscellaneous symbols

type:  print: type:  print:
\hbar  ℏ \ell  ℓ
\imath  ı  \jmath  ȷ
\wp  ℘ \Re  ℜ
\Im  ℑ \partial  ∂
\infty  ∞ \prime  ′
\emptyset ∅ \backslash  \
\forall  ∀ \exists  ∃
\smallint  ∫  \triangle  △
\surd  √ \Vert  ∥
\top  ⊤ \bot  ⊥
\P  ¶ \S  §
\dag  † \ddag  ‡
\flat  ♭ \natural  ♮
\sharp  ♯ \angle  ∠
\clubsuit  ♣ \diamondsuit ♢
\heartsuit ♡ \spadesuit  ♠
\neg  ¬
\Box □ \Diamond ◇
\mho  ℧
type:  print: type:  print:
\hslash  ℏ \complement  ∁
\backprime  ‵  \vartriangle  △
\Bbbk  𝕜 \varnothing  ∅
\diagup  ⧸ \diagdown  ⧹
\blacktriangle  ▲ \blacktriangledown ▼
\triangledown  ▽ \Game  ⅁
\square  □ \blacksquare  ■
\lozenge  ♢ \blacklozenge  ♦
\measuredangle ∡ \sphericalangle  ∢
\circledS  Ⓢ \bigstar  ⋆
\Finv  Ⅎ \eth  ð
\nexists  ∄

This table is divided into two parts. The top part contains the symbols provided by L^{A}TEX; the last three commands require the latexsym package.

The bottom part requires the amssymb package.

Math Spacing Commands

Short form: Full form:  Size: Short form: Full form:
\,  \thinspace  a a \!  \negthinspace
\:  \medspace  a a  \negmedspace
\;  \thickspace a a  \negthickspace
\quad  a a
\qquad  a a

The \medspace, \thickspace, \negmedspace, and \negthickspace commands require the amsmath package.

Delimiters

Name:  Type:  Print: Name:  Type:  Print:
Left paren  (  (  Right paren  )  )
Left bracket  [  [  Right bracket  ]  ]
Left brace  \{  { Right race  \}  }
Reverse slash  \backslash \ Reverse slash  /  /
Left angle  \langle  ⟨  Right angle  \rangle  ⟩
Vertical line  |  |  Double ver.line  \|  ∥
Left floor  \lfloor  ⌊  Right floor  \rfloor  ⌋
Left ceiling  \lceil  ⌈  Right ceiling  \rceil  ⌉
Upper left corner \ulcorner  ⌜ Upper right corner \urcorner ⌝
Lower left corner \llcorner ⌞ Lower right corner \lrcorner ⌟

The corners require the amsmath package.

Name:  type:  print:
Upward arrow  \uparrow  ↑
Double upward arrow  \Uparrow  ⇑
Downward arrow  \downarrow  ↓
Double downward arrow  \Downarrow  ⇓
Up-and-down arrow  \updownarrow  ↕
Double Up-and-down arrow \Updownarrow ⇕

Operators

\arccos  \arcsin  \arctan  \arg
\cos  \cosh  \cot  \coth
\csc  \dim  \exp  \hom
\ker  \lg  \ln  \log
\sec  \sin  \sinh  \tan

\tanh

\varliminf \varlimsup \varinjlim \varprojlim

The \var commands require the amsmath package.

\det \gcd  \inf  \injlim
\lim \liminf \limsup \max
\min \projlim \Pr  \sup

The \injlim and \prolim commands require the amsmath package.

Type:  Inline: Displayed: Type:  Inline: Displayed:
\prod {i=1}ˆ {n}  ∏_{i=1}^{n}  n ; ∏_{i=1}  \coprod {i=1}ˆ {n} ∐_{i=1}^{n}  n ; ∐_{i=1}
\bigcap {i=1}ˆ {n}  ⋂_{i=1}^{n}  n ; ⋂_{i=1}  \bigcup {i=1}ˆ {n}  ⋃_{i=1}^{n}  n ; ⋃_{i=1}
\bigwedge {i=1}ˆ {n} ∧_{i=1}^{n}  ∧_{i=1}^{n}  \bigvee {i=1}ˆ {n}  ⋁_{i=1}^{n}  ∨_{i=1}^{n}
\bigsqcup {i=1}ˆ {n} ⨆_{i=1}^{n}  n ; ⨆_{i=1}  \biguplus {i=1}ˆ {n} ⨄_{i=1}^{n}  n ; ⨄_{i=1}
\bigotimes {i=1}ˆ {n} ⨂_{i=1}^{n}  n ; ⨂_{i=1}  \bigoplus {i=1}ˆ {n} ⨁_{i=1}^{n}  n ; ⨁_{i=1}
\bigodot {i=1}ˆ {n}  ⨀_{i=1}^{n}  ⊙_{i=1}^{n}  \sum {i=1}ˆ {n}  ∑_{i=1}^{n}  n ; ∑_{i=1}

Math Accents

\hat{a}  â \Hat{a}  â \widehat{a} â a\sphat  a
\tilde{a}  ã \Tilde{a} ã \widetilde{a} ã a\sptilde a^{∼}
\bar{a}  ā \Bar{a}  ā
\breve{a}  ă \Breve{a} ă  a\spbreve a˘
\check{a}  ǎ \Check{a} ǎ  a\spcheck a^{∨}
\dot{a}  ȧ \dot{a}  ȧ  a\spdot  a.
\ddot{a}  ä \Ddot{a} ä  a\spddot a..
\dddot\{a\} \begin{matrix}... \\ a \end{matrix}  a\spdddot a...
\ddddot\{a\} \begin{matrix}.... \\ a \end{matrix}
\grave{a}  à \Grave{a} à \imath  ı
\vec{a}  a⃗ \Vec{a}  a⃗ \jmath  ȷ

The \dddot and \ddddot commands and all the captalized commands require the amsmath package; the commands in the

fourth column require the amsxtra package.

Math Font Commands

Type:  Print:
\mathbf{A}  A
\mathit{A}  A
\mathsf{A}  A
\mathrm{A}  A
\mathtt{A}  A
\mathnormal{A} A
\mathbb{A}  𝔸
\mathfrak{A}  𝔄
\mathbf{A}  A
\mathcal{A}  A
\boldsymbol{A}  A
\mathit{\α }  α

The \mathbb, \mathfrak, and \mathcal commands and all the

captalized commands require the amsfonts package. The \boldsymbol command the amssbsy package.